# Fermionic Partial Transpose in the Overlap Matrix Framework for Entanglement Negativity

### v2

@@ Abstract
Over the past two decades, the overlap matrix approach has been developed to compute quantum entanglement in free-fermion systems, particularly to calculate entanglement entropy and entanglement negativity. This method involves the use of partial trace and partial transpose operations within the overlap matrix framework. However, previous studies have only considered the conventional partial transpose in fermionic systems, which does not account for fermionic anticommutation relations. Although the concept of a fermionic partial transpose was introduced in [1], it has not yet been systematically incorporated into the overlap matrix framework. In this paper, we introduce the fermionic partial transpose into the overlap matrix approach, provide a systematic analysis of the validity of partial trace and partial transpose operations, and derive an explicit formula for calculating entanglement negativity in bipartite systems. Additionally, we numerically compute the logarithmic negativity of two lattice models to verify the Gioev-Klich-Widom scaling law. For tripartite geometries, we uncover limitations of the overlap matrix method and demonstrate that the previously reported logarithmic negativity result for a homogeneous one-dimensional chain in a disjoint interval geometry exceeds its theoretical upper bound. Our findings contribute to a deeper understanding of partial trace and partial transpose operations in different representations.
@@ I Introduction
Research on quantum entanglement in recent decades has offered a novel perspective on the analysis of many-body states in condensed matter physics [2, 3, 4, 5]. Unlike classical correlations, quantum entanglement reveals the intrinsic connections between subsystems of a quantum state. This unique property has made quantum entanglement a critical framework for understanding collective behaviors, quantum phase transitions, topological order, and quantum criticality in condensed matter systems [6, 7, 8, 9, 10]. Among the various measures of entanglement, the entanglement entropy has proven to be particularly effective in characterizing bipartite entanglement in the ground states of many-body systems [11, 12, 5].
@@
Compared to other mixed state entanglement measures, entanglement negativity has the advantage of requiring only straightforward linear algebra computations. In this paper, we focus on logarithmic negativity, which is defined as follows. Given the density matrix of a mixed state, such as a reduced density matrix obtained by tracing out region $B$ while retaining $\rho_{A_{1}\cup A_{2}}$, the logarithmic negativity is given by
@@
where $\rho_{A_{1}\cup A_{2}}^{T_{A_{2}}}$ denotes the partial transpose operation on region $A_{2}$ over the reduced density matrix $\rho_{A_{1}\cup A_{2}}$, and the trace norm operation $\text{Tr}\left|O\right|$ represents the sum of the square roots of the eigenvalues of $O^{{\dagger}}O$. If $O$ is Hermitian, the trace norm simplifies to the sum of the absolute eigenvalues of $O$, though this is not always the case. Logarithmic negativity has been extensively applied to investigate entanglement in various many-body systems, including one-dimensional harmonic oscillators [19, 20, 21, 22], spin systems [23, 24, 25, 26, 27, 28, 29, 30], topologically ordered phases [31, 32, 33, 10], and conformal field theory (CFT) [34, 35, 36, 37].
@@
The negativity depends on the definition of partial transpose. In the literature, three distinct forms of partial transposition have been studied: the bosonic partial transpose (bPT), as discussed in Ref. [38, 39]; the untwisted partial transpose (uPT), introduced in Ref. [1, 40, 41, 42]; and the twisted partial transpose (tPT), presented in Ref. [43]. In fermionic systems, the definition of partial transpose must respect the fermionic anticommuting relations. Both uPT and tPT satisfy this requirement, whereas bPT does not. Compared to bPT, uPT and tPT have advantageous properties. For example, it has been shown that a fermionic Gaussian state remains Gaussian after applying uPT or tPT [1, 40, 41, 43], making them convenient for simulating Rényi negativity using the quantum Monte Carlo (QMC) algorithm [42, 44]. Interestingly, the logarithmic negativity calculated using tPT coincides with that of uPT. Therefore, in this work, we focus solely on uPT and omit further discussion of tPT.
@@
However, in practice, the calculation of logarithmic negativity is very challenging, even for free fermion systems. For example, the Green function approach [38, 39, 1, 43] for free-fermion systems is very powerful, but it is still difficult to extract analytical properties of logarithmic negativity. In contrast, the overlap matrix approach has emerged as a more effective tool for computing entanglement entropy and entanglement negativity in free-fermion systems [45, 46, 47, 48, 49]. However, in previous applications of the overlap matrix approach, only bPT was considered. In this paper, we incorporate uPT into the overlap matrix approach. We provide a systematic proof of the validity of the uPT operation, a numerical and analytical confirmation of the partial trace operation in the bipartite geometry of free-fermion systems, and propose analytical results consistent with the CFT method. In the tripartite case, we demonstrate that the previous overlap matrix approach for computing logarithmic negativity exceeds its theoretical upper bound.
@@
This paper is organized as follows. In Sec. II, we introduce the fermionic partial transpose in the overlap matrix approach for entanglement negativity calculation in free fermions system. In Sec. III, we establish the validity of the formula for computing entanglement negativity in bipartite pure states and prove that the partial transpose density matrix in the overlap matrix approach is related to that in Fock space through a similarity transformation. We also calculate two examples, and our numerical results are consistent with the Gioev-Klich-Widom scaling law. In Sec. IV, we study the logarithmic negativity of free fermions in tripartite case, and demonstrate that in mixed states, the equivalence of partial transposition in the overlap matrix approach and in Fock space no longer holds. Consequently, we show that previously reported results for tripartite geometries violate their theoretical upper bound. Finally, in Sec. V, we summarize our findings.
@@ II Overlap matrix approach in free-fermion systems
where the Hermitian matrix $h$ can be diagonalized by a unitary matrix $U$, such that $f_{\alpha}^{\dagger}=\left(\bm{c}^{{\dagger}}U\right)_{\alpha}=\sum_{i=1}^{N}c_{i}^{{\dagger}}U_{i,\alpha}$ is the fermionic operator in the diagonal basis with energy $\epsilon_{\alpha}$, satisfying the anticommutation relation $\left\{f_{\alpha},f_{\beta}^{{\dagger}}\right\}=\delta_{\alpha,\beta}$. Here, $N$ denotes the number of lattice sites. An $M$-particle ground state can be written as
|$\ket{\Psi}=\prod_{\alpha=1}^{M}f_{\alpha}^{{\dagger}}\ket{0},$ (3)
where $M$ lowest energy single particle states are occupied. We consider a bipartite system in which the lattice sites are divided into two regions, labelled by $A$ and $B$. A projection operator $\mathcal{P}_{A\left(B\right)}$ can be defined to only keep the particles in $A\left(B\right)$ for a given state. Specifically, the projection of $f^{{\dagger}}_{\alpha}$ operator satisfies $\mathcal{P}_{A}f_{\alpha}^{\dagger}=\sum_{i\in A}c_{i}^{\dagger}U_{i,\alpha}$ and $\mathcal{P}_{B}f_{\alpha}^{\dagger}=\sum_{i\in B}c_{i}^{\dagger}U_{i,\alpha}$. As expected, the sum of the projections satisfies $\mathcal{P}_{A}f_{\alpha}^{{\dagger}}+\mathcal{P}_{B}f_{\alpha}^{{\dagger}}=f_{\alpha}^{{\dagger}}$, ensuring completeness. An overlap matrix is then introduced
@@
where |$\ket{u_{\alpha}}=f_{\alpha}^{{\dagger}}\ket{0}$ is a single-particle state. It can be easily verified that $M_{A\left(B\right)}$ is Hermitian and $M_{A}+M_{B}=\mathds{1}$. Consequently, the two overlap matrices $M_{A}$ and $M_{B}$ can be simultaneously diagonalized as $\mathcal{U}^{{\dagger}}M_{A}\mathcal{U}=\mathrm{diag}\left(\left\{P_{\gamma}\right\}_{\gamma=1}^{M}\right)$ and $\mathcal{U}^{{\dagger}}M_{B}\mathcal{U}=\mathrm{diag}\left(\left\{1-P_{\gamma}\right\}_{\gamma=1}^{M}\right)$. The eigenvalues $P_{\gamma}$ are restricted to the range $\left[0,1\right]$ (see Appendix A for a proof). We can use $\mathcal{U}$ to define a new basis
@@
|$\ket{\Psi}=\text{e}^{\text{i}\theta}\prod_{\gamma=1}^{M}d_{\gamma}^{\dagger}\ket{0},$ (6)
where the phase factor $\text{e}^{\text{i}\theta}=\det\left(\mathcal{U}^{{\dagger}}\right)$. Another interesting property of this new basis is that the region $A$ and $B$ part operators separate,
@@
where the basis is ordered as {|$\left\{\ket{0_{A\gamma}1_{B\gamma}},\ket{1_{A\gamma}0_{B\gamma}}\right\}$ and the creation operator acts as $d_{A\left(B\right)\gamma}^{{\dagger}}\ket{0}=\ket{1_{A\left(B\right)\gamma}}$. Notably, the original density matrix, which initially has dimensions $2^{N}\times 2^{N}$, is effectively reduced to $2^{M}\times 2^{M}$. This reduction occurs because other rays in the Hilbert space do not contribute to |$\ket{\Psi}$, and all other matrix elements in $\rho$ are strictly zero. It is particularly worth mentioning that the most time-consuming computation in the overlap matrix approach is the diagonalization of the $M\times M$ overlap matrix, whereas in the Green’s function approach, the most time-consuming computation is the diagonalization of the $N\times N$ Green’s function matrix. Therefore, in general, the overlap matrix approach is more advantageous, as $M$ is smaller than $N$.
@@
where the basis is ordered as {|$\left\{\ket{0_{A\gamma}},\ket{1_{A\gamma}}\right\}$. Before performing the partial transposition, it is useful to outline some key properties of the uPT, as discussed in Ref. [41]. The definition in Fock space is given by
(|$\hskip 10.00002pt(|\{n_{j}\}_{A},\{n_{j}\}_{B}\rangle\langle\{\bar{n}_{j}\}_{A},\{\bar{n}_{j}\}_{B}|)^{T_{B}^{f}}$ (11)
@@ III Logarithmic negativity of free fermions system in bipartite case
III.1 Validity of partial transpose in overlap matrix approach
Before discussing the numerical results, we begin with the proof of Eq. (18). For the Hamiltonian in Eq. (2), given that the density matrix of the ground state has a tensor product structure, and using Eq. (15), the logarithmic negativity
Hence, one can analyze each $\rho_{\gamma}^{T_{B}^{f}}$ individually. We prove that the partially transposed density matrix $\rho_{\gamma}$ in the overlap matrix approach is related to its counterpart in Fock space through a unitary transformation.
Figure 1: Bipartite system for a chain and a honeycomb lattice. (a) A contiguous region $A$ consisting of $l$ sites is selected within a system of total size $L$. (b) For the honeycomb lattice, the system is partitioned such that one region forms a corner, with each edge consisting of $l=L/3$ unit cells. The entire system contains $L$ unit cells along each lattice vector, allowing us to examine the scaling law in two dimensions.
Figure 2: Logarithmic negativity $\mathcal{E}$ in different lattice systems. (a) $\mathcal{E}$ as a function of $\ln l$ in a one-dimensional infinite chain for a bipartite system. The total chain length is set to $1000$ sites. Two parameter choices are compared with the analytical formula (black line), both confirming the linear relationship between $\mathcal{E}$ and the logarithm of the subsystem size. (b) $\mathcal{E}/L$ versus $\ln L$ in a honeycomb lattice. The results indicate that logarithmic negativity follows the two-dimensional scaling law $\mathcal{E}\sim L\ln L$. The slope varies for different $\mu/t$, reflecting changes in the Fermi surface.
Proof. We expand the partially transposed single-mode density matrix in the overlap matrix approach ${\rho}_{\gamma}^{T_{B}^{f}}$ into Fock space
${\rho}_{\gamma}^{T_{B}^{f}}$ $=\sum_{i\in A}\sum_{i^{\prime}\in A}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i^{\prime}}c_{i}^{{\dagger}}\ket{0}\bra{0}c_{i^{\prime}}$ (20)
$\hskip 10.00002pt+\sum_{i\in B}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i^{\prime}}c_{i}^{{\dagger}}\ket{0}\bra{0}c_{i^{\prime}}$
$\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i^{\prime}}\ket{0}\bra{0}c_{i^{\prime}}c_{i}$
$\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(U\mathcal{U}\right)_{i^{\prime}\gamma}c_{i}^{{\dagger}}c_{i^{\prime}}^{{\dagger}}\ket{0}\bra{0}.$ | .
$\tilde{\rho}_{\gamma}^{T_{B}^{f}}$ $=\sum_{i\in A}\sum_{i^{\prime}\in A}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i^{\prime}}c_{i}^{{\dagger}}\ket{0}\bra{0}c_{i^{\prime}}$ (21)
$\hskip 10.00002pt+\sum_{i\in B}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i^{\prime}}c_{i}^{{\dagger}}\ket{0}\bra{0}c_{i^{\prime}}$
$\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{\gamma i^{\prime}}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{\gamma i}\ket{0}\bra{0}c_{i^{\prime}}c_{i}$
$\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{i^{\prime}\gamma}c_{i}^{{\dagger}}c_{i^{\prime}}^{{\dagger}}\ket{0}\bra{0},$
[…]
Using Eq. (14), the identity $\mathcal{E}\left(\rho_{\gamma}^{T_{B}^{f}}\right)=\mathcal{E}\left(\tilde{\rho}_{\gamma}^{T_{B}^{f}}\right)$ holds, allowing us to compute bipartite negativity in free-fermion systems via the overlap matrix approach. In Appendix C, we verify the eigenvalues of Eq. (17) through exact diagonalization and obtain consistent results, indicating that a similarity transformation maps $\rho^{T_{B}^{f}}$ to $\tilde{\rho}^{T_{B}^{f}}$. This confirms that the overlap matrix can be effectively used to analyze the negative spectrum[50, 51]. Intriguingly, we find that the eigenvalues of $\rho^{T_{B}^{f}}$ are either purely imaginary or real.
III.2 Two numerical examples
In the following, we consider two free fermion systems. The Hamiltonian we discuss here is
[…]
where $\mu>0$ is the chemical potential, and $\langle i,j\rangle$ denotes the nearest neighbor hopping, with $t>0$ the hopping amplitude.
  1. a.
    One-dimensional infinite chain. Consider an infinite lattice system $\left(L\to\infty\right)$ with periodic boundary conditions (PBC), i.e. $c_{N+1}=c_{1}$. In a bipartite setup, as shown in Fig. 1(a), the overlap matrix takes the form of a Toeplitz matrix.
    $\left[M_{A}\right]_{\alpha,\beta}=L^{-1}\sum_{\nu=0}^{l-1}\text{e}^{\text{i}\frac{2\pi}{L}\nu\left(\beta-\alpha\right)}.$ (23)

### v3

@@ Abstract
Over the past two decades, the overlap matrix approach has been developed to compute quantum entanglement in free-fermion systems, particularly to calculate entanglement entropy and entanglement negativity. This method involves the use of partial trace and partial transpose operations within the overlap matrix framework. However, in previous studies, only the conventional partial transpose in fermionic systems has been considered, which does not account for fermionic anticommutation relations. Although the concept of a fermionic partial transpose was introduced by Shapourian et al. [Phys. Rev. B 95, 165101 (2017)], it has not yet been systematically incorporated into the overlap matrix framework. In this paper, we introduce the fermionic partial transpose into the overlap matrix approach, provide a systematic analysis of the validity of partial trace and partial transpose operations, and derive an explicit formula for calculating entanglement negativity in bipartite systems. Additionally, we numerically compute the logarithmic negativity of two lattice models to verify the Gioev-Klich-Widom scaling law. For tripartite geometries, we uncover limitations of the overlap matrix method and demonstrate that the previously reported logarithmic negativity result for a homogeneous one-dimensional chain in a disjoint interval geometry exceeds its theoretical upper bound.
@@ I Introduction
Research on quantum entanglement in recent decades has offered an additional perspective on the analysis of many-body states in condensed matter physics [1, 2, 3, 4]. Unlike classical correlations, quantum entanglement reveals the intrinsic connections between subsystems of a quantum state. This unique property has made quantum entanglement a critical framework for understanding collective behaviors, quantum phase transitions, topological order, and quantum criticality in condensed matter systems [5, 6, 7, 8, 9]. Among the various measures of entanglement, the entanglement entropy has proven to be particularly effective in characterizing bipartite entanglement in the ground states of many-body systems [10, 11, 4].
@@
Compared with other mixed state entanglement measures, entanglement negativity has the advantage of requiring only straightforward linear algebra computations. In this paper, we focus on logarithmic negativity, which is defined as follows. Given the density matrix of a mixed state, such as a reduced density matrix obtained by tracing out region $B$ while retaining $\rho_{A_{1}\cup A_{2}}$, the logarithmic negativity is given by
@@
where $\rho_{A_{1}\cup A_{2}}^{T_{A_{2}}}$ denotes the partial transpose operation on region $A_{2}$ over the reduced density matrix $\rho_{A_{1}\cup A_{2}}$, and the trace norm operation $\text{Tr}\left|O\right|$ represents the sum of the square roots of the eigenvalues of $O^{{\dagger}}O$. If $O$ is Hermitian, the trace norm simplifies to the sum of the absolute eigenvalues of $O$. Logarithmic negativity has been extensively applied to investigate entanglement in various many-body systems, including one-dimensional harmonic oscillators [18, 19, 20, 21], spin systems [22, 23, 24, 25, 26, 27, 28, 29], topologically ordered phases [30, 31, 32, 9], and conformal field theory (CFT) [33, 34, 35, 36].
@@
The negativity depends on the definition of partial transpose. In the literature, three distinct forms of partial transpose have been studied: bosonic partial transpose (bPT), as discussed in Ref. [37, 38]; untwisted partial transpose (uPT), introduced in Ref. [39, 40, 41, 42]; and twisted partial transpose (tPT), presented in Ref. [43]. In fermionic systems, the definition of partial transpose must respect the fermionic anticommuting relations. Both uPT and tPT satisfy this requirement, whereas bPT does not. Compared with bPT, uPT and tPT have advantageous properties. For example, it has been shown that a fermionic Gaussian state remains Gaussian after applying uPT or tPT [39, 40, 41, 43], making them convenient for simulating Rényi negativity using the quantum Monte Carlo (QMC) algorithm [42, 44]. Interestingly, the logarithmic negativity calculated using tPT coincides with that of uPT. Therefore, in this work, we focus solely on uPT and omit further discussion of tPT.
@@
However, in practice, the calculation of logarithmic negativity is very challenging, even for free fermion systems. For example, the Green function approach [37, 38, 39, 43, 45, 46] for free-fermion systems is very powerful, but it is still difficult to extract analytical properties of logarithmic negativity. In contrast, the overlap matrix approach has emerged as a more effective tool for computing entanglement entropy and entanglement negativity of pure states in free-fermion systems [47, 48, 49, 50, 51, 52]. However, in previous applications of the overlap matrix approach, only bPT was considered. In this paper, we incorporate uPT into the overlap matrix approach. We provide a systematic proof of the validity of the uPT operation, as well as a numerical and analytical confirmation of the partial trace operation in the bipartite geometry of free-fermion systems, and propose analytical results consistent with the CFT method. Additionally, in the tripartite case, we demonstrate that the previous overlap matrix approach for computing logarithmic negativity exceeds its theoretical upper bound.
@@
This paper is organized as follows. In Sec. II, we introduce the fermionic partial transpose in the overlap matrix approach for entanglement negativity calculation in free fermions systems. We then present a rigorous proof that establishes the validity of this framework for bipartite pure states. Specifically, we demonstrate that the partially transposed density matrix from the overlap matrix method is related to its Fock space counterpart through a similarity transformation. In Sec. III, we analyze two illustrative examples, where our numerical results show excellent agreement with the Gioev-Klich-Widom scaling law. In Sec. IV, we study the logarithmic negativity of free fermions in the tripartite case, and demonstrate that in mixed states, the equivalence of partial transpose in the overlap matrix approach and in Fock space no longer holds. Consequently, we show that previously reported results for tripartite geometries violate their theoretical upper bound. Finally, in Sec. V, we summarize our findings.
@@ II Overlap matrix approach in free-fermion systems
where the Hermitian matrix $h$ can be diagonalized by a unitary matrix $U$, such that $f_{\alpha}^{\dagger}=\left(\textbf{c}^{{\dagger}}U\right)_{\alpha}=\sum_{i=1}^{N}c_{i}^{{\dagger}}U_{i,\alpha}$ is the fermionic operator in the diagonal basis with energy $\epsilon_{\alpha}$, satisfying the anticommutation relation $\left\{f_{\alpha},f_{\beta}^{{\dagger}}\right\}=\delta_{\alpha,\beta}$. Here, $N$ denotes the number of lattice sites. An $M$-particle ground state can be written as
[…]
where $M$ lowest energy single particle states are occupied.
II.1 Methodology
We consider a bipartite system in which the lattice sites are divided into two regions, labeled by $A$ and $B$. A projection operator $\mathcal{P}_{A\left(B\right)}$ can be defined to only keep the particles in $A\left(B\right)$ for a given state. Specifically, the projection of the $f^{{\dagger}}_{\alpha}$ operator satisfies $\mathcal{P}_{A}f_{\alpha}^{\dagger}=\sum_{i\in A}c_{i}^{\dagger}U_{i,\alpha}$ and $\mathcal{P}_{B}f_{\alpha}^{\dagger}=\sum_{i\in B}c_{i}^{\dagger}U_{i,\alpha}$. As expected, the sum of the projections satisfies $\mathcal{P}_{A}f_{\alpha}^{{\dagger}}+\mathcal{P}_{B}f_{\alpha}^{{\dagger}}=f_{\alpha}^{{\dagger}}$, ensuring completeness. An overlap matrix is then introduced
@@ II.1 Methodology
[…]
@@
[…]
where the phase factor $\exp\left(\text{i}\theta\right)=\det\left(\mathcal{U}^{{\dagger}}\right)$. Another interesting property of this new basis is that the region $A$ and $B$ part operators separate
@@
[…]
@@
[…]
$\hskip 10.00002pt(|\{n_{j}\}_{A},\{n_{j}\}_{B}\rangle\langle\{\bar{n}_{j}\}_{A},\{\bar{n}_{j}\}_{B}|)^{T_{B}^{f}}$ (11)
@@
Figure 1: Comparison of the eigenvalues of $\rho_{A}$, after reordering and removing the zero components, between the partial trace in the overlap matrix approach and in Fock space. We consider an $N=12$ and $\mu/t=1$ homogeneous one-dimensional chain in two bipartite geometries: (a) an adjacent interval case where the system is evenly divided into regions AA and $B$; (b) a disjoint interval case where the system is evenly partitioned into $A_{1}$, $B_{1}$, $A_{2}$, and $B_{2}$ sequentially [53].
II.2 A general proof
Incredibly when we perform both the partial trace and partial transpose in the new basis, the outcome appears to be identical to that observed in the Fock basis. A rigorous proof is therefore required to validate this equivalence.
Partial trace. The reduced density matrix obtained by analytically performing the partial trace in the Fock basis is compared with the reduced density matrix obtained by the overlap matrix method. The comparison reveals that these two matrices are identical. A detailed analytical derivation confirming this result is provided in Appendix D. We also verify this identical relationship numerically by comparing the eigenvalues of Eq. (10) with exact diagonalization results. We consider two types of bipartite geometries: adjacent intervals and disjoint intervals. As shown in Fig. 1, both geometries yield identical eigenvalues for the partial trace in different representations. Our analytical proof and numerical verifications justify that the entanglement spectrum [54, 55, 56, 57] can be reliably obtained using the overlap matrix approach.
The key reason behind the validity of the partial trace in the overlap matrix approach lies in its semilocal nature, meaning that it remains associated with either region $A$ or $B$. Consequently, tracing out region $A$ or $B$ does not lead to the loss of information about the remaining region, which is different from the Jordan-Wigner transformation [58, 59, 60].
Partial transpose. For the Hamiltonian in Eq. (2), given that the density matrix of the ground state has a tensor product structure, and using Eq. (15), the logarithmic negativity is
Hence, one can analyze each $\rho_{\gamma}^{T_{B}^{f}}$ individually. We prove that the partially transposed density matrix $\rho_{\gamma}$ in the overlap matrix approach is related to its counterpart in Fock space through a unitary transformation. We expand the partially transposed single-mode density matrix in the overlap matrix approach ${\rho}_{\gamma}^{T_{B}^{f}}$ into Fock space
[…]
$|.\displaystyle\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(U\mathcal{U}\right)_{i^{\prime}\gamma}c_{i}^{{\dagger}}c_{i^{\prime}}^{{\dagger}}\ket{0}\bra{0}$.
[…]
$|,\displaystyle\hskip 10.00002pt{-\text{i}}\sum_{i\in A}\sum_{i^{\prime}\in B}\left(U\mathcal{U}\right)_{i\gamma}\left(\mathcal{U}^{{\dagger}}U^{{\dagger}}\right)_{i^{\prime}\gamma}c_{i}^{{\dagger}}c_{i^{\prime}}^{{\dagger}}\ket{0}\bra{0}$,
evidently, transforming $\tilde{\rho}_{\gamma}^{T_{B}^{f}}$ into ${\rho}_{\gamma}^{T_{B}^{f}}$can be achieved through the unitary transformation $c_{i}\to\exp\left[-2\text{i arg}(\phi_{\gamma i)}\right]c_{i}$, provided that $i\in B$, where $\phi_{\gamma i}=\left(U\mathcal{U}\right)_{\gamma i}$.
Using Eq. (14), the identity $\mathcal{E}\left(\rho_{\gamma}^{T_{B}^{f}}\right)=\mathcal{E}\left(\tilde{\rho}_{\gamma}^{T_{B}^{f}}\right)$ holds, allowing us to compute bipartite negativity in free-fermion systems via the overlap matrix approach.
In Appendix C, we verify the eigenvalues of Eq. (17) through exact diagonalization and obtain consistent results, indicating that a similarity transformation maps $\rho^{T_{B}^{f}}$ to $\tilde{\rho}^{T_{B}^{f}}$. This confirms that the overlap matrix can be effectively used to analyze the negativity spectrum [61, 62].
@@ III Logarithmic negativity of free fermions system in bipartite case
Figure 2: Bipartite system for a chain and a honeycomb lattice. (a) A contiguous region AA consisting of $l$ sites is selected within a system of total size $L$. (b) For the honeycomb lattice, the system is partitioned such that one region forms a corner, with each edge consisting of $l=L/3$ unit cells. The entire system contains $L$ unit cells along each lattice vector, allowing us to examine the scaling law in two dimensions.
In this section, we consider two free fermion systems. The Hamiltonian we discuss here is
$H=-\mu\sum_{i}c_{i}^{{\dagger}}c_{i}-t\sum_{\langle i,j\rangle}\left(c_{i}^{{\dagger}}c_{j}+\text{H.c.}\right),$ (22)
where $\mu>0$ is the chemical potential, $\langle i,j\rangle$ denotes the nearest-neighbor hopping, and $t>0$ is the hopping amplitude.
III.1 One-dimensional infinite chain
Consider an infinite lattice system $\left(L\to\infty\right)$ with periodic boundary conditions (PBCs), i.e. $c_{N+1}=c_{1}$. In a bipartite setup, as shown in Fig. 2(a), the overlap matrix takes the form of a Toeplitz matrix.
$\left[M_{A}\right]_{\alpha,\beta}=L^{-1}\sum_{\nu=0}^{l-1}\exp\left[\text{i}\frac{2\pi}{L}\nu\left(\beta-\alpha\right)\right].$ (23)